OpenStudy (anonymous):

If 6x^2 + 5x - 6 = (6x + m)(x - n) for all values of x, then the value of m could be A. -9 B. -4 C. 4 D. 5 E. 6

OpenStudy (anonymous):

First expand (6x+m)(x-n)

OpenStudy (anonymous):

I have no idea.

OpenStudy (anonymous):

Foil

OpenStudy (anonymous):

I'm going to need step-by-step help on this one.

OpenStudy (anonymous):

Do you know FOIL? How about distributive property?

OpenStudy (anonymous):

I've heard about foil, but can't remember what it stands for.

jigglypuff314 (jigglypuff314):

First Outer Inner Last

OpenStudy (anonymous):

My, mom is almost home so I have to go. I probably be here tomorrow. Bye, thanks for the help on the last question.

OpenStudy (anonymous):

*I'll

OpenStudy (anonymous):

I at least became both of you fans.

jigglypuff314 (jigglypuff314):

@wio can you still help? I realized that I actually don't know how to do this either >,<

OpenStudy (anonymous):

Bye sorry that I have to go.

jigglypuff314 (jigglypuff314):

see ya ;)

OpenStudy (anonymous):

Do you want to do the problem jiggly?

jigglypuff314 (jigglypuff314):

yeah :) so 6x^2 - 6xn + mx - mn now what?

OpenStudy (anonymous):

Now put it in the form (...)x^2 + (...)x + (...) However (...) should not contain any x in it.

jigglypuff314 (jigglypuff314):

so 6x^2 - (6n - m)x - mn

OpenStudy (anonymous):

(6)x^2 - (6n - m)x - (mn) = (6)x^2 - (-5)x - (6)

OpenStudy (anonymous):

6=6 6n-m = -5 mn = 6

OpenStudy (anonymous):

To solve for m, we isolate n: n = 6/m Then put it into other equation 6(6/m) - m = -5

OpenStudy (anonymous):

I assume you did foil correctly.

jigglypuff314 (jigglypuff314):

ooohhhhhh!!! yeah so 36 - m^2 = -5m then m^2 - 5m - 36 = 0 (m - 9) (m + 4) = 0 ok, that works! thanks! :D

OpenStudy (anonymous):

Good job guys. You both are so smart, I couldn't have done that.

jigglypuff314 (jigglypuff314):

wio's smarter ;)